OpenStudy (anonymous):

Find the x value for the point that splits segment CD in half if point C is located at (−2, 4) and point D is located at (3, 7).

OpenStudy (zzr0ck3r):

so you need the midpoint of a line segment.

OpenStudy (anonymous):

how would I find that

OpenStudy (zzr0ck3r):

the midpoint of a line created by connecting two points \((a,b),(c,d)\) is \(\\(\frac{a+c}{2},\frac{b+d}{2})\)

OpenStudy (zzr0ck3r):

so \((\frac{-2+3}{2},\frac{4+7}{2})\)